# DeCode: Decoupling Content and Delivery for Medical QA

### v1

@@ 1 Introduction
Large language models (LLMs) have recently achieved strong performance on a variety of medical natural language processing tasks, most notably medical question answering (QA), where models are evaluated on their ability to generate correct responses to clinically relevant questions (singhal2025toward; nori2023can). This progress has been demonstrated across a growing collection of medical QA benchmarks, spanning multiple-choice and generative settings, professional examination-style questions, and open-domain clinical knowledge assessments (jin2021disease; pal2022medmcqa). Collectively, results on these evaluations suggest that contemporary LLMs exhibit substantial medical knowledge and reasoning capability under standardized testing conditions (saab2024capabilities; openai2024o1).
Existing medical QA benchmarks, however, are predominantly designed to measure answer correctness or reasoning accuracy, often via exact-match, multiple-choice selection, or expert-graded factual validity. While these metrics are well-suited for assessing knowledge recall and clinical reasoning, they provide only a partial characterization of model behavior in patient-facing or clinical communication settings (gong2025knowledge; tu2025conversational). In particular, such evaluations do not capture whether model responses are understandable, appropriately calibrated to patient context, or aligned with norms of safe and empathetic medical communication.
This limitation motivates the need for evaluation frameworks that extend beyond accuracy-based metrics. OpenAI HealthBench was introduced to address this gap by evaluating medical LLM outputs along multiple qualitative dimensions, including context seeking, emergency referrals, and responding under uncertainty, in addition to factual correctness (arora2025healthbench). Unlike prior medical QA datasets, which typically assume a single correct answer independent of delivery style or audience, HealthBench explicitly models the interactional aspects of medical responses, enabling a more fine-grained analysis of clinically relevant response quality.
Empirical results on HealthBench further show that models with comparable accuracy on traditional medical QA benchmarks can exhibit substantial variation across other non-accuracy dimensions, revealing a misalignment between standardized QA performance and patient-centered context awareness (arora2025healthbench). Together, these findings suggest that accuracy alone is insufficient as a proxy for real-world clinical readiness and underscore the importance of multidimensional evaluation for medical LLMs.
@@ 2 Related Work
Early evaluations of large language models (LLMs) in medical question answering have primarily focused on standardized multiple-choice benchmarks, including MedQA jin2021disease, MedMCQA pal2022medmcqa, and PubMedQA jin2019pubmedqa. These benchmarks have catalyzed substantial research on assessing and improving medical knowledge in LLMs singhal2025toward; nori2023can; saab2024capabilities; jeong2024improving; li2024agenthospital; wu2025automedprompt. However, such evaluations remain inherently static and accuracy-centric, limiting their ability to assess communicative competence, contextual sensitivity, and patient-centered delivery beyond factual correctness gong2025knowledge.
HealthBench arora2025healthbench introduces a multidimensional evaluation framework for medical QA based on open-ended, multi-turn clinical conversations. Unlike traditional multiple-choice benchmarks, HealthBench employs physician-authored rubrics to assess behavioral dimensions such as clinical accuracy, communication quality, and contextual awareness, enabling a more comprehensive evaluation of medical QA systems beyond factual correctness.
MuSeR zhou2025muser targets HealthBench by proposing a self-refinement training framework in which a student LLM is supervised using high-quality responses from a reference teacher model. The student generates an initial response, performs structured self-assessment across multiple dimensions, and produces a refined final answer. While effective, this approach relies on computationally intensive data synthesis and additional training, limiting its applicability to trainable, open-source LLMs.
In parallel, multi-agent frameworks have been proposed to address complex medical QA by decomposing reasoning across specialized roles. MedAgents tang2024medagents employs role-playing specialists for debate-based hypothesis refinement, while MDAgents kim2024mdagents dynamically configures expert teams based on query complexity. More recent approaches further extend this paradigm: KAMAC wu2025kamac introduces on-demand expert recruitment to address knowledge gaps during generation, and AI Hospital fan2025aihospital evaluates agent-based systems in interactive patient simulation environments. However, these methods primarily emphasize diagnostic reasoning and accuracy on traditional benchmarks, often overlooking how complex reasoning outcomes are translated into clear, user-aligned responses.
@@ 4.1 Dataset and Evaluation
We evaluate on OpenAI HealthBench (arora2025healthbench), which contains 5,0005{,}000 simulated multi-turn patient–clinician conversations ending in a user query. Each conversation is annotated by medical professionals and assigned to one of seven themes: emergency referrals, context seeking, global health, health data tasks, complex responses, hedging, and communication. HealthBench additionally provides physician-authored rubrics per conversation, grouped into five evaluation axes: accuracy, completeness, communication quality, context awareness, and instruction following. For more details regarding the conversation themes and evaluation axes, please refer to the original paper arora2025healthbench.
@@
We follow the official HealthBench protocol (arora2025healthbench): each conversation is graded using its rubric and scored by GPT-4.1 openai2025gpt4.1. Reported numbers are the mean normalized score over the evaluated set.
@@ 4.2 Implementation Details
We use OpenAI o3 openai2025o3 as the primary base model and additionally evaluate GPT-5 openai2025gpt5, Claude-Sonnet 4.5 anthropic2025claude45, and DeepSeek R1 deepseekr1 to assess generalization across model families.
@@
We compare against: (i) Zero-shot, which directly prompts the base LLM to respond from the conversation history; (ii) MDAgents (kim2024mdagents), which adapts team structure and recruits specialized experts based on query complexity, followed by consensus; and (iii) KAMAC (wu2025kamac), which recruits experts on demand during generation. For KAMAC, we let the model choose the initial number of experts and fix the discussion to two rounds.
@@ 5.2 Generalizability Across Backbone Models
GPT-5 OpenAI o3 Claude 4.5 DeepSeek R1
@@
Emergency Ref. 49.8 60.8 (+11.0) 27.0 59.1 (+32.1) 18.5 50.2 (+31.7) 19.9 33.0 (+13.1)
Context Seeking 41.1 57.6 (+16.5) 30.0 58.3 (+28.3) 12.1 46.8 (+34.7) 15.0 32.0 (+17.0)
Global Health 34.1 48.4 (+14.3) 31.8 49.8 (+18.0) 7.3 33.9 (+26.6) 15.6 22.3 (+6.7)
Health Data Tasks 25.8 42.9 (+17.1) 17.0 35.6 (+18.6) 15.0 35.4 (+20.4) 3.1 22.0 (+18.9)
Communications 36.2 51.5 (+15.3) 29.2 43.8 (+14.6) 16.6 39.8 (+23.2) 16.1 18.0 (+1.9)
Hedging 37.4 53.1 (+15.7) 30.9 54.6 (+23.7) 13.1 48.7 (+35.6) 18.3 34.1 (+15.8)
Complex Resp. 31.5 40.3 (+8.8) 24.4 41.3 (+16.9) 15.2 27.5 (+12.3) 14.8 15.9 (+1.1)
@@
Accuracy 46.1 57.2 (+11.1) 45.6 54.3 (+8.7) 28.6 49.5 (+20.9) 30.5 32.6 (+2.1)
Completeness 29.4 56.6 (+27.2) 30.7 58.8 (+28.1) 3.7 44.1 (+40.4) 15.6 27.8 (+12.2)
Comm. Quality 61.0 48.7 (-12.3) 55.5 54.2 (-1.3) 67.3 53.5 (-13.8) 60.9 58.4 (-2.5)
Cont. Awareness 31.0 43.4 (+12.4) 4.0 40.5 (+36.5) 1.5 30.9 (+29.4) 0.0 19.1 (+19.1)
Inst. Following 51.8 43.4 (-8.4) 45.8 46.5 (+0.7) 45.7 43.6 (-2.1) 44.8 42.5 (-2.3)
@@
Based on the zero-shot performance of the base LLMs, health data tasks emerge as a particularly challenging category for GPT-5, OpenAI o3, and DeepSeek R1. In contrast, Claude-4.5 exhibits its weakest performance on global health tasks. Across evaluation axes, context awareness and completeness are especially challenging: OpenAI o3, Claude-4.5, and DeepSeek R1 all record single-digit scores on these dimensions in certain cases, indicating systematic deficiencies in handling complex contextual and informational requirements.
Consistent with the observations in Section 5.1, DeCode delivers substantial improvements over the corresponding zero-shot baselines across all tested models. Notably, Claude-4.5, which attains an initial overall score of $12.4\%$, improves to $40.0\%$ when integrated with DeCode. Similarly, the strongest baseline model, GPT-5, improves from $36.0\%$ to $50.7\%$. Importantly, in all experiments the underlying base LLM remains unchanged. By explicitly decoupling content from delivery, DeCode systematically enhances the performance of diverse base LLMs across nearly all medical QA scenarios. These results demonstrate that the benefits of DeCode are robust and largely LLM-agnostic, extending across a wide range of model architectures and providers.
5.3 Comparison with Multi-Agent Frameworks
Table 3: Comparison with leading multi-agent frameworks. We evaluate representative multi-agent methods, KAMAC wu2025kamac and MDAgents kim2024mdagents, alongside standard zero-shot prompting and DeCode, all using OpenAI o3 on the HealthBench hard subset. Inline deltas are computed relative to the zero-shot baseline; deltas greater than 2 points are highlighted.
Zero-Shot MDAgents KAMAC DeCode
Metric ($\uparrow$, %) (kim2024mdagents) (wu2025kamac) (Ours)
@@ 5.3 Comparison with Multi-Agent Frameworks
Emergency Referrals 27.0 36.3 (+9.3) 33.1 (+6.1) 59.1 (+32.1)
Context Seeking 30.0 34.4 (+4.4) 27.4 (-2.6) 58.3 (+28.3)
Global Health 31.8 43.0 (+11.2) 30.2 (-1.6) 49.8 (+18.0)
Health Data Tasks 17.0 19.1 (+2.1) 7.7 (-9.3) 35.6 (+18.6)
Communications 29.2 38.0 (+8.8) 32.9 (+3.7) 43.8 (+14.6)
Hedging 30.9 39.2 (+8.3) 29.9 (-1.0) 54.6 (+23.7)
Complex Responses 24.4 31.6 (+7.2) 28.5 (+4.1) 41.3 (+16.9)
@@
Accuracy 45.6 52.9 (+7.3) 43.8 (-1.8) 54.3 (+8.7)
Completeness 30.7 45.8 (+15.1) 36.1 (+5.4) 58.8 (+28.1)
Communication Quality 55.5 49.5 (-6.0) 46.7 (-8.8) 54.2 (-1.3)
Context Awareness 4.0 4.6 (+0.6) 0.0 (-4.0) 40.5 (+36.5)
Instruction Following 45.8 53.5 (+7.7) 36.2 (-9.6) 46.5 (+0.7)
In this experiment, we evaluate representative multi-agent medical QA frameworks on the HealthBench hard subset using OpenAI o3 as the base LLM. Specifically, we consider MDAgents kim2024mdagents and KAMAC wu2025kamac, with results summarized in Table 3.
@@
Taken together, these results suggest that when and how experts are introduced plays a critical role in multi-agent medical QA performance. While complexity-driven, fixed-team orchestration promotes stability and coherent reasoning, dynamically expanding agent sets during generation may introduce coordination overhead that outweighs its potential benefits—particularly when all experts are instantiated from the same underlying base LLM.
@@ 6 Conclusion
Future work may explore mechanisms for caching and updating patient-specific information across multi-round interactions, enabling stronger long-term personalization. Additionally, extending the DeCode paradigm beyond medical QA to other high-stakes, user-centered domains represents a promising direction. Taken together, these results suggest that DeCode provides a principled and extensible foundation for advancing contextualized medical question answering.
@@ 7 Limitations
Appendix A Prompt Templates

### v2

@@ 1 Introduction
Large language models (LLMs) have recently achieved strong performance on a variety of medical natural language processing tasks, most notably medical question answering (QA), where models are evaluated on their ability to generate correct responses to clinically relevant questions (Singhal et al., 2025; Nori et al., 2023). This progress has been demonstrated across a growing collection of medical QA benchmarks, spanning multiple-choice and generative settings, professional examination-style questions, and open-domain clinical knowledge assessments (Jin et al., 2021; Pal et al., 2022). Collectively, results on these evaluations suggest that contemporary LLMs exhibit substantial medical knowledge and reasoning capability under standardized testing conditions (Saab et al., 2024; OpenAI, 2024).
Existing medical QA benchmarks, however, are predominantly designed to measure answer correctness or reasoning accuracy, often via exact-match, multiple-choice selection, or expert-graded factual validity. While these metrics are well-suited for assessing knowledge recall and clinical reasoning, they provide only a partial characterization of model behavior in patient-facing or clinical communication settings (Gong et al., 2025; Tu et al., 2025). In particular, such evaluations do not capture whether model responses are understandable, appropriately calibrated to patient context, or aligned with norms of safe and empathetic medical communication.
This limitation motivates the need for evaluation frameworks that extend beyond accuracy-based metrics. OpenAI HealthBench was introduced to address this gap by evaluating medical LLM outputs along multiple qualitative dimensions, including context seeking, emergency referrals, and responding under uncertainty, in addition to factual correctness (Arora et al., 2025). Unlike prior medical QA datasets, which typically assume a single correct answer independent of delivery style or audience, HealthBench explicitly models the interactional aspects of medical responses, enabling a more fine-grained analysis of clinically relevant response quality.
Empirical results on HealthBench further show that models with comparable accuracy on traditional medical QA benchmarks can exhibit substantial variation across other non-accuracy dimensions, revealing a misalignment between standardized QA performance and patient-centered context awareness (Arora et al., 2025). Together, these findings suggest that accuracy alone is insufficient as a proxy for real-world clinical readiness and underscore the importance of multidimensional evaluation for medical LLMs.
@@ 2 Related Work
Early evaluations of large language models (LLMs) in medical question answering have primarily focused on standardized multiple-choice benchmarks, including MedQA Jin et al. (2021), MedMCQA Pal et al. (2022), and PubMedQA Jin et al. (2019). These benchmarks have catalyzed substantial research on assessing and improving medical knowledge in LLMs Singhal et al. (2025); Nori et al. (2023); Saab et al. (2024); Jeong et al. (2024); Li et al. (2024); Wu et al. (2025a). However, such evaluations remain inherently static and accuracy-centric, limiting their ability to assess communicative competence, contextual sensitivity, and patient-centered delivery beyond factual correctness Gong et al. (2025).
HealthBench Arora et al. (2025) introduces a multidimensional evaluation framework for medical QA based on open-ended, multi-turn clinical conversations. Unlike traditional multiple-choice benchmarks, HealthBench employs physician-authored rubrics to assess behavioral dimensions such as clinical accuracy, communication quality, and contextual awareness, enabling a more comprehensive evaluation of medical QA systems beyond factual correctness.
MuSeR Zhou et al. (2025) targets HealthBench by proposing a self-refinement framework in which a student LLM is guided by high-quality responses from a reference teacher model. In its original formulation, MuSeR relies on data synthesis and supervised training: the student generates an initial response, performs structured self-assessment across multiple dimensions, and produces a refined final answer. While this training-based pipeline is computationally intensive and primarily applicable to trainable, open-source LLMs, the core self-refinement procedure can also be applied at inference time, enabling response refinement without model distillation or fine-tuning.
In parallel, multi-agent frameworks have been proposed to address complex medical QA by decomposing reasoning across specialized roles. MedAgents Tang et al. (2024) employs role-playing specialists for debate-based hypothesis refinement, while MDAgents Kim et al. (2024) dynamically configures expert teams based on query complexity. More recent approaches further extend this paradigm: KAMAC Wu et al. (2025b) introduces on-demand expert recruitment to address knowledge gaps during generation, and AI Hospital Fan et al. (2025) evaluates agent-based systems in interactive patient simulation environments. However, these methods primarily emphasize diagnostic reasoning and accuracy on traditional benchmarks, often overlooking how complex reasoning outcomes are translated into clear, user-aligned responses.
@@ 4.1 Dataset and Evaluation
We evaluate on OpenAI HealthBench (Arora et al., 2025), which contains 5,0005{,}000 simulated multi-turn patient–clinician conversations ending in a user query. Each conversation is annotated by medical professionals and assigned to one of seven themes: emergency referrals, context seeking, global health, health data tasks, complex responses, hedging, and communication. HealthBench additionally provides physician-authored rubrics per conversation, grouped into five evaluation axes: accuracy, completeness, communication quality, context awareness, and instruction following. For more details regarding the conversation themes and evaluation axes, please refer to the original paper Arora et al. (2025).
@@
We follow the official HealthBench protocol (Arora et al., 2025): each conversation is graded using its rubric and scored by GPT-4.1 OpenAI (2025a). Reported numbers are the mean normalized score over the evaluated set.
@@ 4.2 Implementation Details
We use OpenAI o3 OpenAI (2025c) as the primary base model and additionally evaluate GPT-5.2 OpenAI (2025b), Claude-Sonnet 4.5 Anthropic (2025), and DeepSeek R1 DeepSeek-AI et al. (2025) to assess generalization across model families.
@@
We compare against: (i) Zero-shot, which directly prompts the base LLM using the conversation history; (ii) MDAgents (Kim et al., 2024), which recruits specialized experts based on query complexity and aggregates their responses; (iii) KAMAC (Wu et al., 2025b), which dynamically recruits experts during generation; and (iv) MuSeR (Zhou et al., 2025), which applies a self-refinement generation strategy. For KAMAC, the model selects the initial number of experts and the discussion is fixed to two rounds. For MuSeR, we use the authors’ self-refinement prompts at inference time, without model distillation or fine-tuning. A detailed cost and latency breakdown for all frameworks is included in the appendix.
@@ 5.2 Generalizability Across Backbone Models
GPT-5.2 OpenAI o3 Claude 4.5 DeepSeek R1
@@
Emergency Ref. 53.6 65.8 (+12.2) 27.0 59.1 (+32.1) 18.5 50.2 (+31.7) 19.9 33.0 (+13.1)
Context Seeking 40.7 62.6 (+21.9) 30.0 58.3 (+28.3) 12.1 46.8 (+34.7) 15.0 32.0 (+17.0)
Global Health 31.9 51.8 (+19.9) 31.8 49.8 (+18.0) 7.3 33.9 (+26.6) 15.6 22.3 (+6.7)
Health Data Tasks 34.0 52.1 (+18.1) 17.0 35.6 (+18.6) 15.0 35.4 (+20.4) 3.1 22.0 (+18.9)
Communications 33.8 55.7 (+21.9) 29.2 43.8 (+14.6) 16.6 39.8 (+23.2) 16.1 18.0 (+1.9)
Hedging 37.8 60.7 (+22.9) 30.9 54.6 (+23.7) 13.1 48.7 (+35.6) 18.3 34.1 (+15.8)
Complex Resp. 35.1 44.0 (+8.9) 24.4 41.3 (+16.9) 15.2 27.5 (+12.3) 14.8 15.9 (+1.1)
@@
Accuracy 49.4 62.6 (+13.2) 45.6 60.9 (+8.7) 28.6 49.5 (+20.9) 30.5 32.6 (+2.1)
Completeness 24.9 56.6 (+31.7) 30.7 58.8 (+28.1) 3.7 44.1 (+40.4) 15.6 27.8 (+12.2)
Comm. Quality 68.0 54.8 (-13.2) 55.5 54.2 (-1.3) 67.3 53.5 (-13.8) 60.9 58.4 (-2.5)
Cont. Awareness 33.0 50.7 (+17.7) 4.0 40.5 (+36.5) 1.5 30.9 (+29.4) 0.0 19.1 (+19.1)
Inst. Following 56.8 48.0 (-8.8) 45.8 46.5 (+0.7) 45.7 43.6 (-2.1) 44.8 42.5 (-2.3)
@@
Based on the zero-shot performance of the base LLMs, health data tasks emerge as a particularly challenging category for GPT-5.2, OpenAI o3, and DeepSeek R1. In contrast, Claude-4.5 exhibits its weakest performance on global health tasks. Across evaluation axes, context awareness and completeness are especially challenging: OpenAI o3, Claude-4.5, and DeepSeek R1 all record single-digit scores on these dimensions in certain cases, indicating systematic deficiencies in handling complex contextual and informational requirements.
Consistent with the observations in Section 5.1, DeCode delivers substantial improvements over the corresponding zero-shot baselines across all tested models. Notably, Claude-4.5, which attains an initial overall score of $12.4\%$, improves to $40.0\%$ when integrated with DeCode. Similarly, the strongest baseline model, GPT-5.2, improves from $36.6\%$ to $56.0\%$. Importantly, in all experiments the underlying base LLM remains unchanged. By explicitly decoupling content from delivery, DeCode systematically enhances the performance of diverse base LLMs across nearly all medical QA scenarios. These results demonstrate that the benefits of DeCode are robust and largely LLM-agnostic, extending across a wide range of model architectures and providers.
5.3 Comparison with LLM Frameworks
Table 3: Comparison with leading LLM-based frameworks. We evaluate MDAgents Kim et al. (2024), KAMAC Wu et al. (2025b), and MuSeR Zhou et al. (2025), together with standard zero-shot prompting and DeCode, all using OpenAI o3 on the HealthBench hard subset. Inline deltas are reported relative to the zero-shot baseline, with improvements greater than 2 points highlighted.
Zero-Shot MDAgents KAMAC MuSeR DeCode
Metric ($\uparrow$, %) (Kim et al., 2024) (Wu et al., 2025b) (Zhou et al., 2025) (Ours)
@@ 5.3 Comparison with LLM Frameworks
Emergency Referrals 27.0 36.3 (+9.3) 33.1 (+6.1) 53.3 (+26.3) 59.1 (+32.1)
Context Seeking 30.0 34.4 (+4.4) 27.4 (-2.6) 60.1 (+30.1) 58.3 (+28.3)
Global Health 31.8 43.0 (+11.2) 30.2 (-1.6) 44.4 (+12.6) 49.8 (+18.0)
Health Data Tasks 17.0 19.1 (+2.1) 7.7 (-9.3) 37.2 (+20.2) 35.6 (+18.6)
Communications 29.2 38.0 (+8.8) 32.9 (+3.7) 40.4 (+11.2) 43.8 (+14.6)
Hedging 30.9 39.2 (+8.3) 29.9 (-1.0) 51.2 (+20.3) 54.6 (+23.7)
Complex Responses 24.4 31.6 (+7.2) 28.5 (+4.1) 37.7 (+13.3) 41.3 (+16.9)
@@
Accuracy 45.6 52.9 (+7.3) 43.8 (-1.8) 53.5 (+7.9) 54.3 (+8.7)
Completeness 30.7 45.8 (+15.1) 36.1 (+5.4) 49.1 (+18.4) 58.8 (+28.1)
Communication Quality 55.5 49.5 (-6.0) 46.7 (-8.8) 54.1 (-1.4) 54.2 (-1.3)
Context Awareness 4.0 4.6 (+0.6) 0.0 (-4.0) 46.0 (+42.0) 40.5 (+36.5)
Instruction Following 45.8 53.5 (+7.7) 36.2 (-9.6) 46.2 (+0.4) 46.5 (+0.7)
In this experiment, we evaluate representative LLM-based medical QA frameworks on the HealthBench hard subset using OpenAI o3 as the base LLM. Specifically, we consider MDAgents Kim et al. (2024), KAMAC Wu et al. (2025b), and MuSeR Zhou et al. (2025) with results summarized in Table 3.
@@
Among the compared LLM-based frameworks, MuSeR’s self-refinement approach yields the largest performance gains, particularly on context seeking and health data tasks, and achieves the strongest improvement in context awareness. This is consistent with MuSeR’s design, where structured self-assessment guides targeted response refinement to recover missing contextual signals. Our results show that such self-refinement remains effective even when applied purely at inference time. However, MuSeR does not explicitly disentangle clinical content from discourse strategy, limiting its ability to independently optimize medical accuracy and communicative quality—an aspect directly addressed by DeCode.
Taken together, these results indicate that multi-agent medicalQA performance depends critically on how contextual information and communicative intent are structured during generation. While fixed-team orchestration and self-refinement improve context sensitivity, dynamically introducing new experts can incur coordination overhead that degrades coherence. Explicitly separating clinical content from discourse strategy provides a more stable alternative, enabling balanced optimization of medical accuracy and communication quality, as demonstrated by DeCode.
@@ 6 Conclusion
Future work may explore mechanisms for caching and updating patient-specific information across multi-round interactions. Additionally, extending the DeCode paradigm beyond medical QA to other user-centered domains represents a promising direction. Taken together, these results suggest that DeCode provides a principled foundation for advancing contextualized medical QA.
@@ 7 Limitations
Appendix A Cost & Latency Analysis
In this section, we analyze the cost and latency of the DeCode framework. In addition, we include the cost and latency of MDAgents Kim et al. (2024), KAMAC Wu et al. (2025b), and MuSeR Zhou et al. (2025). To ensure fair comparison, we assume an output token rate of $70$ tokens/second to avoid latency fluctuations due to API re-routing. The details of the DeCode framework is listed in Table  5. Comparisons between the various LLM-based medical QA frameworks are provided in Table 6.
Table 5: Token usage and latency breakdown of the DeCode framework. We report average input tokens, output tokens, and estimated latency for each DeCode module. Latency is computed assuming a standardized throughput of 70 tokens/s. Since the Profiler and Formulator modules execute in parallel, their combined latency is attributed to the Formulator module.
Component Avg. Input Tok. Avg. Output Tok. Avg. Latency (s)
Background (Profiler) 426.71 229.51 3.29
User Needs (Profiler) 360.71 213.31 3.04
Clinical Indicators (Formulator) 515.71 876.54 12.53
Discourse Strategy (Strategist) 1,420.37 848.34 12.11
Response (Synthesizer) 1,502.16 1,457.63 20.83
DeCode Total (parallelized) 4,225.66 3,625.33 45.47
Table 6: Comparison of computational cost and latency across methods. We report average input tokens, output tokens, estimated cost, and latency per sample. Latency is computed assuming a standardized generation throughput of 70 tokens/s. Bold denotes the best-performing method in each column (lowest cost or lowest latency).
Method Avg. Input Tok. Avg. Output Tok. Avg. Cost (\$) Avg. Latency (s)
MDAgents 44,444 4,909 0.128 70.13
KAMAC 574,438.50 29,610.58 1.386 423.01
MuSeR 6,205.93 2437.13 0.040 48.96
DeCode (Ours) 4,225.66 3,625.33 0.037 45.47
Appendix B Prompt Templates